[figure: 9.7 cm]

TRANSFER-MATRIX SIMULATIONS OF LINEAR MAGNETIC SYSTEMS

R. Matysiak Computational Physics Division, Institute of Physics A. Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland

The quantum transfer-matrix method (QTM) has been applied to study the finitetemperature static properties of the spin S = 1 antiferromagnetic Heisenberg chains in a wide range of the single-ion anisotropy and temperatures. The high-resolution QTM simulation data are obtained for the zero-field susceptibility, specific heat as well as for the field-dependent magnetization. The microscopic parameters of a number of real quasi-one-dimensional compounds are found from fitting procedures, some theoretical approaches are numerically verified and low-temperature extensions of high-temperature series are given in terms of Chebyshev polynomials. An extension of the technique to a non-uniform bound alternating molecular magnets is also put forward.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : R. Matysiak

Address for correspondence : Computational Physics Division, Institute od Physics A. Mickiewicz University, ul. Umultowska 85, 61-614 Poznan, Poland

Email address : matysiak@pearl.amu.edu.pl